OpenStudy (anonymous):

Determine the empirical formula of an oxide of iron which has 69.9% iron and 30.1% dioxygen by mass.

OpenStudy (aaronq):

Fe=69.9g/Fe's molar mass O2=30.1g/O2's molar mass find lowest common factor

OpenStudy (anonymous):

ok i did that......see what to consider O2's molar mass as? It is 16 or 32....thats the major confusion.....the solution booklet says it 2:3 and i got it as 1:1 so plz help @aaronq

OpenStudy (aaronq):

i'm sure you know "di-" means two, so it would be 32

OpenStudy (anonymous):

ya i did that and hence i got 1:1....solve and tell me whether it is 1:1 or 2:3......@aaronq

OpenStudy (aaronq):

69.9/55.85 = 1.25 30.1/32 = 1 FeO2

OpenStudy (aaronq):

because the bottom one counts for 2 oxygens

OpenStudy (aaronq):

Though if you use only one oxygen 69.9/ 55.84/mol = 1.25 30.1/ 15.99/mol = 1.88 then 1.88/1.25=1.5 1 Fe = 1.5O then multiply by 2 to make it a whole number Fe2O3

OpenStudy (anonymous):

@aaronq So wht shud i consider 1 oxygen or 2? The question says diatomic so i am considering 2. What acc to u should be the correct alternative?

OpenStudy (anonymous):

@terenzreignz Hey wat do u say abt this question?

terenzreignz (terenzreignz):

I say... I stay in the Maths section for a reason :)

OpenStudy (anonymous):

@terenzreignz all right no problem.........i think i got my answer...

OpenStudy (dean.shyy):

This may help: http://is.gd/5KzdIo